OpenStudy (princesssleelee):

Let f(x) = 3x2 – x + 2 and g(x) = 5x2 – 1. Find f(g(x)). Show each step of your work.

OpenStudy (priyar):

substitute g(x) in the place of x in the f(x) expression to find f(g(x))...

OpenStudy (princesssleelee):

idk how

OpenStudy (princesssleelee):

put g(x) into the x in the f(x) equation. 5x^2 -1 --\ \textgreater \ 3(5x^2-1)^2 - (5x^2 - 1)^2 Then, all you have to do is simplify. f(g[x]) = 3(25x^4 - 10x^2 + 1) - 25x^2-10x^2+1 f(g[x]) = 3(25x^4 - 10x^2 + 1) - 1(25x^2-10x^2+1) f(g[x]) = 2(25x^4 - 10x^2 + 1) f(g[x]) = 50x^4 - 20x^2 + 2 ????